OpenStudy (anonymous):

right triangle abc. b=90 deg a=60 deg side bc=17.3cm can you find side ac and side ab. use 30-60-90

OpenStudy (anonymous):

Tan 60 = BC/AB ===> AB=BC/Tan(60) =9.99 cm AC= Rac((AB^2)+(BC^2)) ===> AC=19.98 cm

OpenStudy (anonymous):

So What's the answer please

OpenStudy (anonymous):

AB= 9.99 cm and AC=19.98 cm

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

look, we have the same numbers! haha.

OpenStudy (anonymous):

Ab=10 ac=20